# Probing nuclear structure with the Balitsky-Kovchegov equation in full impact-parameter dependence

### v1

@@ 1 Introduction
The work is organised as follows: Sec. 2 presents an overview of the BK formalism and its extension to nuclear targets, together with a description of the approach to modelling the tetrahedral oxygen target. In Sec. 4, we present the predictions for the nuclear structure functions and the vector meson production cross-sections and show the comparison to data, and Sec. 5 briefly summarises the presented work.
@@ 2 Formalism overview
$N(r,b,\theta,\eta=0)=1-\exp\left(-\frac{1}{4}(Q_{\rm s0}0^{2}r^{2})^{\gamma}T_{\rm p}(b,r)(1+c\cos(2\theta))\right)$ (4)
@@
Table 1: Parameters of the solution to the BK including explicit dependence on ($r,b,\theta$) according to [28]. Parameters were fitted for quark masses $m_{\rm uds}=0.1~\mathrm{GeV}$, $m_{\rm c}=1.27~\mathrm{GeV}$ and $m_{b}=4.18~\mathrm{GeV}$.
@@ 3 BK equation solutions
Figure 3: Comparison of the dipole amplitudes at the initial condition (full) and evolved to rapidity 5 (dashed) for a proton target (teal) and lead nucleus (ochre). The shown amplitudes are central (i.e. at small impact parameter $b=0.01~\mathrm{GeV}^{-1}$) and at $\theta=0$ for the standard BK equation $\kappa=1$ (left panel) and its linearised version $\kappa=0$ (right panel).
@@ Acknowledgements
This work was partially funded by the Czech Science Foundation (GAČR), project No. 22-27262S.

### v2

@@ 1 Introduction
The work is organised as follows: Sec. 2 presents an overview of the BK formalism and its extension to nuclear targets, together with a description of the approach to modelling the tetrahedral oxygen target. In Sec. 3, we compare the dipole amplitudes obtained from the standard and linearised BK equation for proton and lead targets, Sec. 4 presents the predictions for the nuclear structure functions and the vector meson production cross-sections and show the comparison to data, and Sec. 5 briefly summarises the presented work.
@@ 2 Formalism overview
$N(r,b,\theta,\eta=0)=1-\exp\left(-\frac{1}{4}(Q_{\rm s0}^{2}r^{2})^{\gamma}T_{\rm p}(b,r)(1+c\cos(2\theta))\right)$ (4)
@@
Due to the computational complexity of the 3D BK equation, the parameters in Table 1 do not come from a full fitting procedure, but rather from an estimate based on a coarse scanning of the parameter space. The palette of experimental measurements used for this estimate consists of proton structure functions [8], differential and total cross-sections of vector-meson production ($\phi$ [7, 42], $\rho$ [20, 77, 33, 43, 6], $\omega$ [50, 35], $J/\psi$ [7, 42, 19, 3, 10, 17, 41], $\psi(2S)$ [3, 1], and $\Upsilon(1S)$ [2, 45, 34, 44, 16]) and DVCS [4], all of them in virtuality range $Q^{2}\in[0,35]~{\rm GeV}^{2}$, resulting in $\chi^{2}/ndof=6.4$. While this $\chi^{2}$ may seem large, it should be noted that this study is simultaneously taking into account a large variety of experimental observables rather than a narrow set of data for one particular process. Given that several of the included data sets have very narrow experimental uncertainties, while the present framework is expected to carry at least a few-percent systematic theory uncertainty (e.g., missing higher-order effects), a $\chi^{2}$ close to unity is not expected; substantially smaller values would likely indicate that theory uncertainties are being effectively absorbed into the fit rather than genuinely predicted. Furthermore, no K-factor has been introduced to improve the data description.
Table 1: Parameters of the solution to the BK including explicit dependence on ($r,b,\theta$) according to [40]. Parameters were obtained for quark masses $m_{\rm uds}=0.1~\mathrm{GeV}$, $m_{\rm c}=1.27~\mathrm{GeV}$ and $m_{b}=4.18~\mathrm{GeV}$.
@@ 3 BK equation solutions
Figure 3: Comparison of the dipole amplitudes at the initial condition (full) and evolved to rapidity 4 (dashed) and 8 (dotted) for a proton (teal) and lead (ochre) target. The shown amplitudes are central (i.e. at small impact parameter $b=0.01~\mathrm{GeV}^{-1}$) and at $\theta=0$ for the standard BK equation $\kappa=1$ (left panel) and its linearised version $\kappa=0$ (right panel).
Note that while, in the particular phase-space point shown in Fig. 3, the peak value of the dipole amplitude decreases between the initial condition and $\eta=4$, the physically relevant evolution takes place at smaller values of $r$, to the left of the peak. Values of $r\geq 1/\Lambda_{\text{QCD}}$ correspond to the non-perturbative region, and this part of phase-space is suppressed by the convolution with (mostly perturbatively calculated) wave-functions in calculations of measurable observables. Here, the amplitude increases steadily, and the evolution resembles the standard travelling wave” known from solutions of the only-$r$-dependent BK equation. Furthermore, it is this low-$r$ region that has the dominant effect on the observables such as the structure functions, which consistently increase with rapidity, despite the decrease of the amplitude peak. This feature is natural to impact-parameter-dependent solutions of the BK equation and has been reported in Ref. [29].
@@ Acknowledgements
This work was partially funded by the Czech Science Foundation (GAČR), project No. 22-27262S. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254). M. Matas acknowledges the support by the CTU Mobility Project MSCA-F-CZ-III (reg.no. CZ.02.01.01/00/22_010/0008601).